rohani buys a washing machine by paying a RM 400 deposit and paying the balance RM 1772 after 5 months. If the interest is 10% compounded monthly, calculate the cash value of the washing machine.


1
Expert's answer
2021-08-11T04:50:05-0400

Solution:

Let the balance after deposit be RM "x" .

Then, after 5 months, amount"=x(1+\dfrac rn)^{nt}"

"1772=x(1+\dfrac{0.1}{12})^{\frac 5{12}\times12}
\\ 1772=x(1.042366)
\\ x=1699.97\approx1700"

Thus, the cash value of the washing machine is RM (1700+400) = RM 2100.